OpenStudy (anonymous):

find a formula for the inverse of the function f(x)=sqr 10-3x

OpenStudy (anonymous):

square both sides : f(x)^2 = 10 - 3x 3x = 10 - f(x)^2 x = (10-f(x)^2)/3 let f-1(x) be the inverse of f(x) so : f-1(x) = (10-x^2)/3

OpenStudy (anonymous):

f(x)^2 where is from

OpenStudy (anonymous):

sqaure both sides i wrote it in the first line

OpenStudy (anonymous):

the question is f(x)= SQR root 10-3x

OpenStudy (anonymous):

right .. and now if you square both sides (raise to the power of 2) you get f(x)^2 = 10 -3x

OpenStudy (anonymous):

f(x)10-3x

OpenStudy (anonymous):

\[f(x) = \sqrt{10-3x}\] square both sides and get : \[{f(x)}^2 = 10-3x\]

OpenStudy (anonymous):

i thing some is wrong

OpenStudy (anonymous):

any way thanks for help

OpenStudy (anonymous):

nothing is wrong .. sqrt(a) = b if you square both sides you get a = b^2 same as sqrt(25) = 5 square both sides you get 25 = 5^2 25 = 25

OpenStudy (anonymous):

i know but the question is f(x)=10-3x Sqr

OpenStudy (anonymous):

\[f(x) = \sqrt{10-3x}\]

OpenStudy (anonymous):

like this ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[f(x) = {(10-3x)}^2\] like this ?

OpenStudy (anonymous):

find formula for the inverse of the function f(x)=